Suppose a firm’s inverse demand curve is given by P=120-.5Q and its cost equation is C=420+60Q+Q2.

  1. Find the firm’s optimal quantity, price and profit (1) by using the profit and marginal profit equation and (2) by setting MR equal to MC. Also provide a graph of MR and MC.
  2. Suppose instead that the firm can sell any and all of its output at the fixed market price P=120. Find the firm’s optimal output.
Answer #1

of output. MC 920 llo Den and MR 60=MC quantity b) P= 120 Pe 120 in dubstituting != 120 P = 10 - 0.5 g .: 120= 120 - 0.58 Q=0